# Riemannian Optimization for Distance Geometry: A Study of Convergence, Robustness, and Incoherence

### v1

@@ Abstract
The problem of recovering a configuration of points from partial pairwise distances, referred to as the Euclidean Distance Geometry (EDG) problem, arises in a broad range of applications, including sensor network localization, molecular conformation, and manifold learning. In this paper, we propose a Riemannian optimization framework for solving the EDG problem by formulating it as a low-rank matrix completion task over the space of positive semi-definite Gram matrices. The available distance measurements are encoded as expansion coefficients in a non-orthogonal basis, and optimization over the Gram matrix implicitly enforces geometric consistency through the triangle inequality, a structure inherited from classical multidimensional scaling. Under a Bernoulli sampling model for observed distances, we prove that Riemannian gradient descent on the manifold of $r$ matrices locally converges linearly with high probability when the sampling probability satisfies $p\geq\mathcal{O}(\nu^{2}r^{2}\log(n)/n)$, where $\nu$ is an EDG-specific incoherence parameter. Furthermore, we provide an initialization candidate using a one-step hard thresholding procedure that yields convergence, provided the sampling probability satisfies $p\geq\mathcal{O}(\nu r^{3/2}\log^{3/4}(n)/n^{1/4})$. A key technical contribution of this work is the analysis of a symmetric linear operator arising from a dual basis expansion in the non-orthogonal basis, which requires a novel application of the Hanson–Wright inequality to establish an optimal restricted isometry property in the presence of coupled terms. Empirical evaluations on synthetic data demonstrate that our algorithm achieves competitive performance relative to state-of-the-art methods. Moreover, we propose a novel notion of matrix incoherence tailored to the EDG setting and provide robustness guarantees for our method.
@@
Euclidean Distance Geometry, Riemannian Optimization, Matrix Completion, Sensor Localization
@@ 1 Introduction
The rapid advancement of technology across various scientific fields has greatly simplified data collection. In many practical applications, however, there are limitations to measurements that can lead to incomplete data. This can be caused by geographic, climatic, or other factors that determine whether a measurement between two points can be obtained, and as such some data may be missing[1, 2]. For instance, in protein structure prediction, nuclear magnetic resonance (NMR) spectroscopy experiments yield spectra for protons that are close together, resulting in incomplete known distance information[3]. Similarly, in sensor networks, we may have mobile nodes with known distances only from fixed anchors [4, 5]. In these and other scenarios, the fundamental problem is determining the configuration of points based on partial information about inter-point distances. This problem is known as the Euclidean distance geometry (EDG) problem, which has numerous applications throughout the applied sciences [6, 7, 8, 9, 10, 11, 12, 13, 14, 15].
@@
We note that $\bm{X}\bm{1}=\bm{0}$ also implies that $\bm{U}^{\top}\bm{1}=\bm{0}$. In many practical scenarios, the distance matrix may be incomplete, making classical MDS inapplicable for determining the point configuration. However, notice that $\mathrm{rank}(\bm{X})\leq r$, and one can show that $\mathrm{rank}(\bm{D})\leq r+2$ [20]. This implies that when $r\ll n$, which is often the case in practice, $\bm{X}$ and $\bm{D}$ are low-rank matrices. This allows us to utilize a rich library of tools from low-rank matrix completion. With this in mind, one technique is to directly apply matrix completion techniques on $\bm{D}$[21]. Let $\Omega\subset\{(i,j)\mid 1\leq i<j\leq n\}$ denote the set of sampled indices corresponding to the strictly upper-triangular part of the distance matrix. Note that, since a distance matrix is hollow and symmetric, it suffices to consider the samples in the upper-triangular part; that is, if $D_{ij}$ is sampled, $D_{ji}$ is also assumed to be sampled. A matrix completion approach would consider the following optimization program to recover $\bm{D}$:
@@
where $\|\cdot\|_{*}=\sum_{i}\sigma_{i}$ denotes the nuclear norm, which serves as a convex surrogate for rank[22]. The main idea of these tools is that, under some assumptions, the nuclear norm minimization program reconstructs the true low-rank squared distance matrix exactly with high probability from $\mathcal{O}(nr\log^{2}(n))$ randomly sampled entries [23, 24, 25, 26, 27]. Another set of techniques [28, 29] focus on recovering the point configuration by using the Gram matrix as an optimization variable, and using only partial information from the entries in $\bm{D}$. Specifically, these works consider the following optimization program for the EDG problem:
$\begin{split}\operatorname*{\mathrm{minimize}}_{\bm{X}\in\mathbb{R}^{n\times n},\,\bm{X}=\bm{X}^{\top},\,\bm{X}\succeq\bm{0},\bm{X}\bm{1}=\bm{0}}&\quad\|\bm{X}\|_{\ast}\\ \end{split}$ (5)
@@
We note that theoretical guarantees for (5) have been established in [31, 28], but still suffer from the lack of scalability of convex techniques. A non-convex Lagrangian formulation was also proposed in [28], yielding strong numerical results but lacking local convergence guarantees. The work in [32] uses a Riemannian manifold approach to develop a conjugate gradient algorithm for estimating the underlying Gram matrix. The theoretical analysis therein shows that the squared distance matrix iterates globally converge to the true squared distance matrix at the sampled entries under three assumptions. However, the relationship between the problem parameters, such as the sampling scheme and sampled entries, and the third assumption remains unclear, as noted in Remark III.8 of the paper. In [30], the authors introduce a Riemannian conjugate gradient method with line search for the EDG problem. The paper provides a local convergence analysis for the case where the entries of the distance matrix are sampled according to the Bernoulli model given a suitable initialization. The initialization method used is known as rank reduction, which begins with initial points embedded in a higher-dimensional space than the target dimension. While [30] demonstrates strong empirical results for this initialization via tests on synthetic data for sensor localization, there are no provable guarantees provided for the initialization. In [33], an asymmetric projected gradient algorithm is proposed that adapts the pseudogradient of an earlier version of this work, seen in [34], using a Burer-Monteiro factorization. Recovery guarantees are provided, but the recovery is established without reference to the restricted isometry of a random measurement operator, indicating a difference in approach to convergence guarantees. Furthermore, the rate of convergence is sublinear in the attractive basin near the solution, and the recovery guarantees scale quadratically with respect to standard incoherence in the matrix completion literature, rather than the EDG-specific incoherence described in detail in this work.The work in [35] proposes a non-convex algorithm for the EDG problem based on the reweighted least squares framework. It considers the case where distance entries are observed uniformly at random and establishes that with $O(\nu r\log(n)$ distance entries, where $\nu$ is the incoherence parameter (see Section 3 for the definition of a weaker form of incoherence used in this paper), are sufficient for local convergence to the ground Gram matrix. However, [35] does not provide a provable initialization scheme or robustness guarantees for the proposed algorithm. We note that the analysis in [35] achieves optimal sample complexity, matching the lower bound established in [36]. However, their results rely on a stronger incoherence condition than ours. In fact, under our milder incoherence assumption, their sample complexity aligns with ours up to constant factors.
@@ 1.1 Contributions
    Algorithmic Framework: We propose a novel non-convex iterative algorithm for the Euclidean Distance Geometry (EDG) problem based on Riemannian optimization. The algorithm performs first-order updates on the manifold of fixed-rank matrices and enjoys low per-iteration computational complexity.
@@
  3. 3.
    Convergence guarantees, sample complexity requirements, and robustness guarantees: We provide rigorous analysis establishing high-probability local convergence of the proposed algorithm to the ground truth configuration with near optimal sample complexity. We also derive sample complexity bounds ensuring that the initialization lies within the basin of attraction and provide robustness guarantees under bounded noise perturbations.
@@
    Novel Analysis and Interpretability: We leverage statistical tools not common in the EDG literature to analyze the local behavior of the algorithm, including a restricted isometry property for a symmetric operator with coupled structure. Furthermore, we offer a new interpretation of matrix incoherence tailored to the EDG setting.
To the best of our knowledge, this is the first non-convex algorithm for the EDG problem that provides provable initialization, provable convergence guarantees, robustness guarantees under noise, and a geometric interpretation of incoherence in the EDG context.
@@ 1.2 Notation
The notation used in this paper is summarized in Table 1. We note that this table describes is what is generally used throughout this paper, but not every assignment is a hard and fast rule. For example, lowercase boldface, such as $\bm{x}$, is denoted as reserved for vectors; however, we extensively use the notation $\bm{w}_{\bm{\alpha}}$ for certain matrices. If there is any contradiction with Table 1, the notation should be clear from context.
@@
$\bm{e}_{i}$ Standard basis vector: 111 at $i$-th position, zeros elsewhere
[…]
@@
$\bm{Y}=\bm{U}\bm{D}\bm{U}^{\top}$ Thin spectral decomposition of symmetric $r$ matrix
@@
$\lambda_{1}(\bm{X})\geq\cdots\geq\lambda_{r}(\bm{X})$ Ordered non-zero eigenvalues of $r$ matrix $\bm{X}$, $\bm{X}$ sometimes omitted
@@
$\bm{x}_{i}$, $\bm{x}^{i}$ $i$-th row and $i$-th column of $\bm{X}$, respectively
@@
$\mathcal{N}_{r}$ Manifold of $r$ matrices
@@ 1.3 Organization
The organization of this paper is as follows. In Section 2, we discuss the requisite background information necessary to understand the work done in this paper. This consists of a brief discussion of low-rank matrix completion and a discussion of EDG, with further background on dual bases and first-order Riemannian methods found in Appendix G. Section 3 gives a detailed discussion of the EDG-specific incoherence condition. Section 4 is a discussion of our proposed methodology for solving the EDG problem using geometric low-rank matrix completion ideas in the developed dual basis framework. Section 5 discusses the underlying assumptions, convergence analysis, initialization guarantees, and robustness results of the proposed algorithm, with most proofs deferred to the Appendices. The convergence analysis leverages the discussed dual basis structure, with properties proven in Appendix A, to get local convergence guarantees, discussed in more detail in Appendices B and C. We additionally provide initialization and robustness guarantees in this section, with relevant proofs in Appendices D and E. Section 7 discusses related geometric approaches in matrix completion, relevant work done in EDG, and a more detailed discussion of geometric approaches to EDG. Section 8 discusses the numerical results of this algorithm, and compares its efficacy to another algorithm in the literature. We conclude the paper in Section 9 with a brief discussion of the work and possible future research directions.
@@ 2.1 Matrix Completion
One of the primary components this work relies on is the field of low-rank matrix completion, where a subset of the entries of a low-rank ground truth matrix $\bm{X}$ are observed. Consider $\bm{X}$ as an $n\times n$ matrix for simplicity, with $\Omega\subset[n]\times[n]$ representing the set of observed indices. Here, a sampling operator $\mathcal{P}_{\Omega}:\mathbb{R}^{n\times n}\to\mathbb{R}^{n\times n}$ is introduced, which aggregates the observed entries of $\bm{X}$ projected onto specific basis elements $\bm{e}_{ij}$:
@@
As minimizing the rank directly is generally a challenging problem [25, 37], relaxations of this problem are often considered. For details on the complexity class of rank constrained problems, we refer the reader to [38]. Exact recovery of $\bm{X}$ from $\mathcal{P}_{\Omega}(\bm{X})$ using a convex relaxation to the nuclear norm, such as the objective described in (4), is a well-studied problem [24, 39, 40] with strong convergence guarantees. This problem is at the core of matrix completion literature, and has inspired work in the completion of distance matrices [29, 28]. However, solving the convex problem is expensive for large matrices, which has led to the consideration of non-convex methodologies to solve the underlying problem. One approach that has received a great deal of attention is the Burer-Monteiro factorization approach, pioneered for semi-definite methods in [41], whereby a low rank matrix $\bm{X}\in\mathbb{R}^{n\times n}$ can be factored into a product $\bm{X}=\bm{A}\bm{B}^{\top}$ for $\bm{A},\bm{B}\in\mathbb{R}^{n\times r}$. Minimizing $\|\mathcal{P}_{\Omega}(\bm{X})-\mathcal{P}_{\Omega}(\bm{A}\bm{B}^{\top})\|_{\mathrm{F}}^{2}$ is a common approach, and is often dealt with using alternating minimization methods in both the noiseless and noisy case [42, 43, 44, 45].
One of the main statistical approaches to analyzing matrix completion problems is through studying the behavior of the sampling operator $\mathcal{P}_{\Omega}$ restricted to a feasible space for recovery. This is formalized by defining, for a $r$ ground truth matrix $\bm{X}$, the tangent space $\mathbb{T}$ at $\bm{X}$ on $\mathcal{N}_{r}$, the manifold of $r$ matrices. Explicitly, we have that
@@
for some constant $c>0$ and some small $\varepsilon_{0}>0$, which depends on both the number of samples and intrinsic properties of the ground truth matrix $\bm{X}$[39]. This property is known as the Restricted Isometry Property (RIP), and variants of this property have been critical to low-rank matrix completion and compressive sensing literature[46].
2.2 Dual Basis Approach to EDG
In the EDG problem, using the relation (2), we can relate each entry of the squared distance matrix to the Gram matrix as follows: $D_{ij}=X_{ii}+X_{jj}-X_{ij}-X_{ji}$. We describe here briefly the dual basis approach introduced in [28]. Given $\bm{\alpha}=(\alpha_{1},\alpha_{2}),\alpha_{1}<\alpha_{2}$, we define the matrix $\bm{w}_{\bm{\alpha}}$ as follows:

### v2

# Provable Non-Convex Euclidean Distance Matrix Completion: Geometry, Reconstruction, and Robustness

@@ Abstract
The problem of recovering the configuration of points from their partial pairwise distances, referred to as the Euclidean Distance Matrix Completion (EDMC) problem, arises in a broad range of applications, including sensor network localization, molecular conformation, and manifold learning. In this paper, we propose a Riemannian optimization framework for solving the EDMC problem by formulating it as a low-rank matrix completion task over the space of positive semi-definite Gram matrices. The available distance measurements are encoded as expansion coefficients in a non-orthogonal basis, and optimization over the Gram matrix implicitly enforces geometric consistency through nonnegativity and the triangle inequality, a structure inherited from classical multidimensional scaling. Under a Bernoulli sampling model for observed distances, we prove that Riemannian gradient descent on the manifold of rank-$r$ matrices locally converges linearly with high probability when the sampling probability satisfies $p\geq\mathcal{O}(\nu^{2}r^{2}\log(n)/n)$, where $\nu$ is an EDMC-specific incoherence parameter. Furthermore, we provide an initialization candidate using a one-step hard thresholding procedure that yields convergence, provided the sampling probability satisfies $p\geq\mathcal{O}(\nu r^{3/2}\log^{3/4}(n)/n^{1/4})$. A key technical contribution of this work is the analysis of a symmetric linear operator arising from a dual basis expansion in the non-orthogonal basis, which requires a novel application of the Hanson-Wright inequality to establish an optimal restricted isometry property in the presence of coupled terms. Empirical evaluations on synthetic data demonstrate that our algorithm achieves competitive performance relative to state-of-the-art methods. Moreover, we provide a geometric interpretation of matrix incoherence tailored to the EDMC setting and provide robustness guarantees for our method.
@@
Euclidean Distance Matrix, Matrix Completion, Riemannian Optimization
@@ 1 Introduction
The rapid advancement of technology across various scientific fields has greatly simplified data collection. In many practical applications, however, there are limitations to measurements that can lead to incomplete data. This can be caused by geographic, climatic, or other factors that determine whether a measurement between two points can be obtained, and as such some data may be missing[1, 2]. For instance, in protein structure prediction, nuclear magnetic resonance (NMR) spectroscopy experiments yield spectra for protons that are close together, resulting in incomplete known distance information[3]. Similarly, in sensor networks, we may have mobile nodes with known distances only from fixed anchors [4, 5]. In these and other scenarios, the fundamental problem is determining the configuration of points based on partial information about inter-point distances. This problem is known as the Euclidean Distance Geometry problem, which has numerous applications throughout the applied sciences [6, 7, 8, 9, 10, 11, 12, 13, 14, 15].
@@
We note that $\bm{X}\bm{1}=\bm{0}$ also implies that $\bm{U}^{\top}\bm{1}=\bm{0}$. In many practical scenarios, the distance matrix may be incomplete, making classical MDS inapplicable for determining the point configuration. However, notice that $\mathrm{rank}(\bm{X})\leq r$, and one can show that $\mathrm{rank}(\bm{D})\leq r+2$ [20]. This implies that when $r\ll n$, which is often the case in practice (e.g., $r=2$ or $3$ in EDMC), $\bm{X}$ and $\bm{D}$ are low-rank matrices. This allows us to utilize a rich library of tools from low-rank matrix completion, and moves us to consider the problem of Euclidean Distance Matrix Completion (EDMC). With this in mind, one technique is to directly apply matrix completion techniques on $\bm{D}$[21]. Let $\Omega\subset\{(i,j)\mid 1\leq i<j\leq n\}$ denote the set of sampled indices corresponding to the strictly upper-triangular part of the distance matrix. Note that, since a distance matrix is hollow and symmetric, it suffices to consider the samples in the upper-triangular part; that is, if $D_{ij}$ is sampled, $D_{ji}$ is also assumed to be sampled. A matrix completion approach would consider the following optimization program to recover $\bm{D}$:
@@
where $\|\cdot\|_{*}=\sum_{i}\sigma_{i}$ denotes the nuclear norm, which serves as a convex surrogate for rank[22]. The main idea of these tools is that, under some assumptions, the nuclear norm minimization program reconstructs the true low-rank squared distance matrix exactly with high probability from $\mathcal{O}(nr\log^{2}(n))$ randomly sampled entries [23, 24, 25, 26, 27]. Another set of techniques [28, 29] focus on recovering the point configuration by using the Gram matrix as an optimization variable, and using only partial information from the entries in $\bm{D}$. Specifically, these works consider the following optimization program for the EDMC problem:
$\begin{split}\operatorname*{\mathrm{minimize}}_{\bm{X}\in\mathbb{R}^{n\times n},\,\bm{X}=\bm{X}^{\top},\,\bm{X}\succeq\bm{0},\bm{X}\bm{1}=\bm{0}}&\quad\|\bm{X}\|_{\ast},\end{split}$ (5)
@@
We note that theoretical guarantees for (5) have been established in [31, 28], but still suffer from the lack of scalability of convex techniques. A non-convex Lagrangian formulation was also proposed in [28], yielding strong numerical results but lacking local convergence guarantees. The work in [32] uses a Riemannian manifold approach to develop a conjugate gradient algorithm for estimating the underlying Gram matrix. The theoretical analysis therein shows that the squared distance matrix iterates globally converge to the true squared distance matrix at the sampled entries under three assumptions. However, the relationship between the problem parameters, such as the sampling scheme and sampled entries, and the third assumption remains unclear, as noted in Remark III.8 of the paper. In [30], the authors introduce a Riemannian conjugate gradient method with line search for the EDMC problem. The paper provides a local convergence analysis for the case where the entries of the distance matrix are sampled according to the Bernoulli model given a suitable initialization. The initialization method used is known as rank reduction, which begins with initial points embedded in a higher-dimensional space than the target dimension. While [30] demonstrates strong empirical results for this initialization via tests on synthetic data for sensor localization, there are no provable guarantees provided for the initialization.
The work by [33] is thematically closer to ours, as it also considers a similar initialization, and proves linear convergence to the ground truth. Therein, the core approach is based on a Burer-Monteiro factorization of the Gram matrix, and this leads to a program where the optimization variable is in terms of points. To solve the resulting optimization problem, a simple gradient descent algorithm with a line search followed by a projection is proposed. The projection is done to ensure incoherence, but requires as input an incoherence parameter and maximum singular value of the true Gram matrix. We highlight few differences of our approach to this work. First, their empirical evaluations assume known incoherence and maximum singular value. We believe such sharp estimates may not be necessary, and the impact of not having this information is not clear. We also note that we provide robustness guarantees in contrast to [33].
The work in [34] proposes a non-convex algorithm for the EDMC problem based on the reweighted least squares framework. It considers the case where distance entries are observed uniformly at random and establishes that with $O(\nu r\log(n)$ distance entries, where $\nu$ is the incoherence parameter (see Section 3 for the definition of a weaker form of incoherence used in this paper), are sufficient for local convergence to the ground Gram matrix. However, [34] does not provide a provable initialization scheme or robustness guarantees for the proposed algorithm. We note that the analysis in [34] achieves optimal sample complexity, matching the lower bound established in [35]. However, their results rely on a stronger incoherence condition than ours. In fact, under our milder incoherence assumption, their sample complexity aligns with ours up to constant factors.
@@ 1.1 Contributions
    Geometric Interpretation of EDMC Coherence: We provide a geometric interpretation of coherence within the specific context of the EDMC problem. We derive both lower and upper bounds for this parameter and discuss the geometries that achieve these bounds. Under a random model of the underlying points, we show that the coherence scales logarithmically, which aligns with the scaling of standard coherence measures.
@@
    Algorithmic Framework: We propose a novel non-convex iterative algorithm for the Euclidean Distance Matrix Completion (EDMC) problem based on Riemannian optimization. The algorithm performs first-order updates on the manifold of fixed-rank matrices and enjoys low per-iteration computational complexity.
  3. 3.
@@
    Convergence guarantees, sample complexity requirements, and robustness guarantees: We provide rigorous analysis establishing high-probability local convergence of the proposed algorithm to the ground truth configuration with near optimal sample complexity. We also derive sample complexity bounds to ensure that the initialization lies within the basin of attraction and to provide robustness guarantees against bounded noise perturbations of the underlying point cloud.
  5. 5.
    Novel Analysis We leverage statistical tools not common in the EDMC literature to analyze the local behavior of the algorithm, including a restricted isometry property for a symmetric operator with coupled structure.
To the best of our knowledge, this is the first non-convex algorithm for the EDMC problem that provides provable initialization, provable convergence guarantees, robustness guarantees under noise, and a geometric interpretation of incoherence in the EDMC context.
@@ 1.2 Notation
The notation used in this paper is summarized in Table 1. This table provides a general description of the conventions used throughout this paper, but not every assignment is a strict rule. For example, lowercase boldface, such as $\bm{x}$, is denoted as reserved for vectors; however, we extensively use the notation $\bm{w}_{\bm{\alpha}}$ for certain matrices. If there is any contradiction with Table 1, the notation should be clear from context.
@@
$\bm{e}_{i}$ Standard basis vector: $1$ at $i$-th position, zeros elsewhere
$\bm{e}_{ij}$ Standard matrix basis: $1$ at $(i,j)$, zeros elsewhere
@@
$\bm{Y}=\bm{U}\bm{D}\bm{U}^{\top}$ Thin spectral decomposition of symmetric rank-$r$ matrix
@@
$\lambda_{1}(\bm{X})\geq\cdots\geq\lambda_{r}(\bm{X})$ Ordered non-zero eigenvalues of rank-$r$ matrix, when clear
from context, $\bm{X}$ is omitted
@@
$\bm{x}_{i}$, $\bm{x}^{i}$ $i$-th row and $i$-th column of $\bm{X}$, respectively,
represented as column vectors.
@@
$\mathcal{N}_{r}$ Manifold of rank-$r$ matrices
@@ 1.3 Organization
The organization of this paper is as follows. In Section 2, we discuss the requisite background information necessary to understand the work done in this paper. This consists of a brief discussion of low-rank matrix completion and a discussion of EDMC, with further background on dual bases and first-order Riemannian methods found in Appendix G. Section 3 gives a detailed discussion of the EDMC-specific incoherence condition. Section 4 is a discussion of our proposed methodology for solving the EDMC problem using geometric low-rank matrix completion ideas in the developed dual basis framework. Section 5 discusses the underlying assumptions, convergence analysis, initialization guarantees, and robustness results of the proposed algorithm, with most proofs deferred to the Appendices. The convergence analysis leverages the discussed dual basis structure, with properties proven in Appendix A, to get local convergence guarantees, discussed in more detail in Appendices B and C. We additionally provide initialization and robustness guarantees in this section, with relevant proofs in Appendices D and E. Section 7 discusses related geometric approaches in matrix completion, relevant work done in EDMC, and a more detailed discussion of geometric approaches to EDMC. Section 8 discusses the numerical results of this algorithm, and compares its efficacy to another algorithm in the literature. We conclude the paper in Section 9 with a brief discussion of the work and possible future research directions.
@@ 2.1 Matrix Completion
This work is related to low-rank matrix completion, where a subset of the entries of a low-rank ground truth matrix $\bm{X}$ are observed. Consider $\bm{X}$ as an $n\times n$ matrix for simplicity, with $\Omega\subset[n]\times[n]$ representing the set of observed indices. Here, a sampling operator $\mathcal{P}_{\Omega}:\mathbb{R}^{n\times n}\to\mathbb{R}^{n\times n}$ is introduced, which aggregates the observed entries of $\bm{X}$ projected onto specific basis elements $\bm{e}_{ij}$:
@@
As minimizing the rank directly is generally a challenging problem [25, 36], relaxations of this problem are often considered. For details on the complexity class of rank constrained problems, we refer the reader to [37]. Exact recovery of $\bm{X}$ from $\mathcal{P}_{\Omega}(\bm{X})$ using a convex relaxation to the nuclear norm, such as the objective described in (4), is a well-studied problem [24, 38, 39]. This problem is at the core of matrix completion literature, and has inspired work in the completion of distance matrices [29, 28]. However, solving the convex problem is expensive for large matrices, which has led to the consideration of non-convex methodologies to solve the underlying problem. One approach that has received a great deal of attention is the Burer-Monteiro factorization approach, pioneered for semi-definite methods in [40], whereby a low rank matrix $\bm{X}\in\mathbb{R}^{n\times n}$ can be factored into a product $\bm{X}=\bm{A}\bm{B}^{\top}$ for $\bm{A},\bm{B}\in\mathbb{R}^{n\times r}$. Minimizing $\|\mathcal{P}_{\Omega}(\bm{X})-\mathcal{P}_{\Omega}(\bm{A}\bm{B}^{\top})\|_{\mathrm{F}}^{2}$ is a common approach, and is often dealt with using alternating minimization methods in both the noiseless and noisy case [41, 42, 43, 44]. Beyond standard matrix completion, these methods have also been applied to other structured problems [45, 46, 47].
One of the main statistical approaches to analyzing matrix completion problems is through studying the behavior of the sampling operator $\mathcal{P}_{\Omega}$ restricted to a feasible space for recovery. This is formalized by defining, for a rank-$r$ ground truth matrix $\bm{X}$, the tangent space $\mathbb{T}$ at $\bm{X}$ on $\mathcal{N}_{r}$, the manifold of rank-$r$ matrices. Explicitly, we have that
@@
for some constant $c>0$ and some small $\varepsilon_{0}>0$, which depends on both the number of samples and intrinsic properties of the ground truth matrix $\bm{X}$[38]. This property is known as the Restricted Isometry Property (RIP), and variants of this property have been critical to low-rank matrix completion and compressive sensing literature[48, 49].
2.2 Dual Basis Approach to EDMC
In the EDMC problem, using the relation (2), we can relate each entry of the squared distance matrix to the Gram matrix as follows: $D_{ij}=X_{ii}+X_{jj}-X_{ij}-X_{ji}$. We describe here briefly the dual basis approach introduced in [28]. Given $\bm{\alpha}=(\alpha_{1},\alpha_{2}),\alpha_{1}<\alpha_{2}$, we define the matrix $\bm{w}_{\bm{\alpha}}$ as follows: